Determine the poles and the residue at each pole of the function f(z)=1/(z-1)2(z+2)


1
Expert's answer
2021-09-02T07:25:28-0400

 If "f(z)" has a pole of order "k" at "z=z_0" then


"rez(f,z_0)=\lim\limits_{z\to z_0}\big(\dfrac{1}{(k-1)!}\dfrac{d^{k-1}}{dz^{k-1}}((z-z_0)^kf(z))\big)"

"z=-2," simple pole

"rez(f,-2)=\lim\limits_{z\to-2}((z+2)f(z))"

"=\lim\limits_{z\to-2}((z+2)(\dfrac{1}{(z+2)(z-1)^2}))"

"=\lim\limits_{z\to-2}(\dfrac{1}{(z-1)^2})=\dfrac{1}{(-2-1)^2}=\dfrac{1}{9}"

"z=1," order "2" pole

"rez(f,1)=\lim\limits_{z\to 1}\big(\dfrac{1}{(2-1)!}\dfrac{d^{2-1}}{dz^{2-1}}((z-1)^2f(z))\big)"

"=\lim\limits_{z\to 1}\big(\dfrac{d}{dz}((z-1)^2(\dfrac{1}{(z+2)(z-1)^2}))\big)"

"=\lim\limits_{z\to 1}\big(\dfrac{d}{dz}(\dfrac{1}{z+2})\big)"

"=\lim\limits_{z\to 1}\big(-\dfrac{1}{(z+2)^2}\big)"

"=-\dfrac{1}{(1+2)^2}=-\dfrac{1}{9}"


"z=-2," simple pole


"rez(f,-2)=\dfrac{1}{9}"



"z=1," order "2" pole


"rez(f,1)=-\dfrac{1}{9}"